### v1

@@ Abstract
Generative AI is a directional technology: it excels at some task combinations and performs poorly at others. Knowledge work is also directional and endogenous: workers can satisfy their job requirements with different combinations of tasks. Studying AI adoption by knowledge workers hence requires comparing two vectors. We develop a high-dimensional model of task choice and technology adoption, with otherwise standard neoclassical assumptions. AI is adopted when its direction is aligned with what the worker values at the margin—the worker’s shadow prices, rather than with what the worker actually does—their activity vector. This yields a cone of adoption that widens as AI capability grows; near the entry threshold, small improvements in capability translate into large expansions in the set of adopted directions. Adoption also has a structured intensive margin: a tool can be worth using but not worth using all the time, generating a region of stable hybrid production between an entry threshold and an all-in threshold. We also show how to derive shadow prices as explicit functions of observable skill and requirement vectors. The framework explains rapid adoption in aligned occupations, heterogeneous adoption elsewhere, and weak correlation with one-dimensional skill measures: the key heterogeneity is directional alignment, not skill level.
@@ 1  Introduction
Generative artificial intelligence (GenAI) based on large language models (Vaswani et al., 2017) shares features of general-purpose technologies (Eloundou et al., 2023) but differs from earlier automation in two ways. First, individual workers can adopt GenAI quickly in knowledge work: the interface is general-purpose and the main inputs are time and attention rather than specialized capital. Adoption can therefore spread through informal and unregulated channels inside organizations. [^1] On worker-driven adoption (”shadow IT”), see Klotz et al. (2019); Raković et al. (2020); for rapid early diffusion, see UBS (2023). Second, capabilities improve quickly but unevenly across tasks. [^2] For software engineering benchmarks, see SWE-bench (Jimenez et al., 2024). The unevenness is described as a ”jagged technological frontier” by Dell’Acqua et al. (2023) and quantified more broadly by Hendrycks et al. (2025). Taken together, these features make it harder to predict who adopts AI and who gains from it.
Measured productivity effects are mixed and do not map cleanly to any single notion of skill. In some settings, usage is already pervasive and mediated by tools that translate intent into working output; in others, adoption is slower and gains depend on complementary investments by firms. [^3] Evidence on productivity effects is mixed across settings and designs; see, e.g., Dell’Acqua et al. (2023); Benzell and Myers (2026) and firm-level evidence on complementarity in Aldasoro et al. (2026). Existing frameworks often compress heterogeneity into task or worker groups or a scalar exposure index, which can obscure the direction-by-direction nature of adoption and motivate a multidimensional model.
To address this, we develop a model where both jobs and AI systems are explicitly multidimensional. A job is a bundle of tasks (Autor, 2013, 2025), and a worker chooses her task mix to maximize output subject to a time or attention budget. Skills shape the feasible set of task combinations, while job requirements shape the value of tasks in production. We model AI as a directional technology that produces tasks in fixed proportions. Using AI requires supervision—time the worker cannot spend on her own production. Because supervision draws on the same scarce budget, adoption is a team-production problem: the combined feasible set is a weighted sum of the worker’s own capabilities and those of AI. When their capabilities point in different directions, the joint production set expands beyond what either achieves in isolation. [^4] The logic extends to AI that can flexibly combine a small number of task directions in a high-dimensional space; see Kaushik et al. (2025). The key analytical step is to evaluate a technology direction at the worker’s marginal valuation: at the optimum, shadow prices convert any task direction into a payoff-relevant scalar.
The model yields a simple organizing prediction: adoption is governed by alignment between the AI direction and what the worker values at the margin. AI is adopted when its direction is sufficiently aligned with the worker’s shadow prices, which are in general different from the worker’s observed activity mix.
We make six contributions.
  1. 1.
    Explain nonlinear adoption without frictions. The model generates a sharp capability threshold for adoption even in a frictionless neoclassical setting. Near this threshold, small improvements in AI capability produce large expansions in the set of adopted tasks. This provides a clean economic explanation for why usage can suddenly take off without invoking hype, network effects, or fixed costs.
  2. 2.
    Separate absolute from comparative advantage. Technological superiority is necessary but not sufficient: AI is used only when its strengths align with what is scarce at the margin in a given worker–job match. This explains why powerful tools are sometimes ignored and why adoption varies widely across firms and roles.
  3. 3.
    Reframe AI as team production. Because AI requires supervision and consumes time and attention, it competes for the same scarce resource as human work. Adoption is therefore a collaboration decision evaluated at shadow prices, not a simple capital–labor substitution story. This shifts the discussion from replacement to directed co-production.
  4. 4.
    Introduce a structured intensive margin. There are two adoption thresholds in our model: an entry threshold and an all-in threshold. Between them, optimal AI use is partial, generating stable hybrid production rather than immediate automation.
  5. 5.
    Establish the worker–job match as the unit of analysis. Adoption depends on the interaction between worker skills and job requirements, not on workers or occupations in isolation. Shadow prices are match-specific, so asking which jobs or which workers are at risk is incomplete without specifying the pairing.
  6. 6.
    Break the scalar exposure paradigm. Adoption cannot be summarized by a one-dimensional skill or exposure index. The relevant sufficient statistic is a vector of shadow prices reflecting multidimensional alignment. As a result, common exposure rankings may systematically misclassify who adopts and who benefits.
These results imply that scalar exposure rankings are generally not sufficient statistics for adoption.
@@
To make the mechanism transparent, our method is intentionally neoclassical. We study a worker doing a given job and abstract from governance and principal-agent frictions, so the worker maximizes output. We characterize autarky task choice using a trade-style duality argument (Dixit and Norman, 1980), which yields shadow (autarky) prices even though tasks are not traded on any market. We then introduce AI as a directional technology that competes for the same scarce budget and derive adoption and intensity conditions in terms of capability and alignment.
For interpretation and measurement, we provide a CES-CET parametric specialization that expresses shadow prices as explicit functions of skill and requirement vectors. The key parameters are the elasticity of substitution across tasks in production ($\sigma$) and the elasticity of transformation in resource use ($\gamma$). When $\gamma\to 0$, workers are rigid specialists and skills alone determine autarky prices; when $\gamma\to\infty$, workers are perfectly flexible and job requirements alone matter. Symmetric logic applies to $\sigma$: low values make tasks strong complements, high values make them close substitutes. In general, both worker skills and job requirements shape autarky prices, consistent with within-job task reallocation (Spitz-Oener, 2006).
The framework also has limitations that suggest where the economics of GenAI must ultimately go. By design, we hold the worker-AI match fixed and treat adoption as an individual optimization problem. We do not model sorting—which tools a worker selects and which workers a tool is designed for—or how organizations structure human-AI teams. In practice, firms set policies, provide complementary investments, and may restrict or encourage usage (Aldasoro et al., 2026). We also treat the technology as exogenous and focus on the worker side of adoption rather than on how training data and market incentives shape the direction of AI capabilities. These abstractions isolate a basic mechanism—directional alignment at shadow prices—that can be embedded in richer models of organizational choice, market structure, and dynamic capability change.
@@
The remainder of the paper proceeds as follows. Section 2 develops the model of task choice and directional technology adoption, derives the adoption condition and the cone of adoption, and characterizes the intensive margin. Section 3 provides a CES-CET parametric specification that yields closed-form expressions for shadow prices and adoption thresholds. Section 4 positions the framework relative to prior task-based models and discusses interpretation, limitations, and connections to the emerging evidence. The Appendix collects proofs and the parametric derivations.
@@ 2  A Model of Digital Technology Adoption
We study a worker doing a given job. We do not model labor market, governance, or principal-agent frictions; consequently, the worker maximizes output. This is the outcome of a richer model in which both employer and worker can observe and contract on task execution. We take the worker-job match as fixed and reserve the study of sorting to future work.
The worker chooses how much to do of each of $N$ tasks. Denote the task quantities by the vector $x\in\mathbb{R}_{+}^{N}$. The set of task bundles the worker can complete in one workday (or other unit of scarce resource) is given by the set $X_{A}$, characterized by its resource use function $g$,
@@
Here $B$ is the amount of the scarce resource (“budget”) the worker has available. We interpret the budget as time or attention, not money. This is a key feature of digital technology adoption: the binding constraint is typically the worker’s cognitive capacity and hours, not financial cost. The set $X_{A}$ is called the Production Possibilities Set (PPS).
@@ 2.1  Technology expands the production possibilities set
To use AI, the worker has to use some of the same resources that constrain them in their own work. For example, instructing AI agents and reviewing their work takes time, and this limits the amount of AI that can be used in a workday. [^6] Time as a binding constraint is central to the economics of digital goods. Goodman et al. (2026) leverage Becker’s time allocation theory to study consumer demand for time-intensive platforms like Facebook and Instagram, finding that time shares and time costs significantly influence substitution patterns across online activities. We assume AI is competing for the same resource budget $B$, so if the worker spends a fraction $\lambda$ of their budget working with AI, only $(1-\lambda)B$ is left for doing other tasks. [^7] For simplicity, we assume there is a linear transformation between resources spent on AI supervision and resources spent on human work. This is eminently true if the resource is time: one hour lost on an activity is one hour gained on another. A more general formulation could model the production possibilities set as constrained by several resources, such as time, attention, mental energy, and AI may not use them in the same proportion as human tasks. We conjecture our results would survive as long as AI is sufficiently substitutable to human tasks in the resource constraint. We denote the capability of technology with $\chi\geq 0$, which acts as a productivity shifter. A unit of technology $\|t\|$ costs $1/\chi$ units of the resource $B$. Continuing with the AI example, more capable AI tools can get more done per hour of human supervision.
@@
Absolute advantage is necessary but not sufficient for technology adoption. A counterexample is illustrated in Figure 4. The technology already has an absolute advantage, so the worker could use it to expand their PPS. In their current job, however, their activities are directionally far from the technology, so it is not worth adopting it. [^9] Geometrically, until the technology vector pierces the budget set given by the hyperplane separate the PPS and the production function upper contour set, its convex hull with the PPS is also strictly within the budget set, extending no revenue increasing opportunities to the worker.
@@ 2.3  The intensive margin of technology
Suppose $F$ is strictly concave and $C^{2}$. Let $y=B\chi t$ and $X_{T}=\mathrm{conv}(X_{A}\cup\{y\})$. Then:
@@
Uniqueness of $\tilde{x}_{*}$ follows from strict concavity of $F$ over the convex set $X_{T}$. The decomposition follows from the convex hull representation (Proposition 2.3). Uniqueness of $\lambda_{*}$ follows from strict concavity of $f(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$, and $x_{*}$ is then pinned down by $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$. Uniqueness of $p_{100}$ follows from $C^{2}$ differentiability of $F$. See Appendix A.1. ∎
@@
Suppose $F$ is strictly concave, $C^{2}$, and homogeneous of degree one. Let $\lambda_{*}(\chi)$ denote the unique optimal technology intensity from Proposition 2.7. Then:
@@
Cases 1–2 follow from the boundary conditions of the univariate function $f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]$. For case 3, the strict inequality $\chi_{100}>\chi_{0}$ follows from strict concavity: the supporting hyperplanes at $x_{A}$ and at the technology direction $t$ are distinct whenever $t$ is not collinear with $x_{A}$. See Appendix A.2. ∎
@@ 2.4  More than one technology
    All-in: The worker goes all-in on technology $K+1$ if $\chi_{K+1}>\varrho_{K}(p_{100,K+1})/p_{100,K+1}^{\prime}t_{K+1}$ , and $\chi_{K+1}/\chi_{k}>p_{100,K+1}^{\prime}t_{k}^{\prime}p_{100,K+1}^{\prime}t_{1}$ for all k=1,2,,Kk=1,2,...,K, where $p_{100,K+1}=\nabla F(t_{K+1})/\|\nabla F(t_{K+1})\|$.
@@ 3.1  CES job requirements and CET skills
Skills (CET). The resource use function is constant elasticity of transformation (CET):
@@ 3.2  Autarky allocation in closed form
$x_{i}^{A}=B\cdot\frac{s_{i}^{\gamma/(\gamma+\sigma)}\,\theta_{i}^{\sigma/(\gamma+\sigma)}}{\Phi^{\sigma/(\gamma+\sigma)}},\qquad i=1,\ldots,N,$
@@
The allocation formula shows that the worker concentrates activity where skills and requirements overlap. The exponents make the role of curvature transparent. When $\sigma$ is high, tasks are easier to substitute in production, so the worker tilts more strongly toward high-requirement activities (higher weight on $\theta_{i}$). When $\gamma$ is high, it is easier to reallocate effort across tasks inside the constraint set, so differences in skills translate more strongly into differences in chosen activity levels.

### v2

@@ Abstract
Generative AI is directional: it performs well in some task directions and poorly in others. Knowledge work is directional and endogenous as well: workers can satisfy the same job requirements with different mixes of tasks. We develop a high-dimensional model of AI adoption in which a worker uses a tool when it raises their output. Both the worker and the AI tool can perform a variety of tasks, which we model as convex production possibility sets. Because the tool requires supervision from the worker’s own time and attention budget, adoption is a team-production decision, similar to hiring a coworker. The key sufficient statistics are the worker’s pre-AI shadow prices: these equal the output gain from a small relaxation in each task direction, and they generally differ from the worker’s observed activity mix. As AI capability improves, the set of adopted directions expands in a cone centered on these autarky prices. Near the entry threshold, small capability improvements generate large extensive-margin expansions in adoption. The model also delivers a structured intensive margin: between the entry and all-in thresholds, optimal use is partial. We parametrize the model in a simple but flexible way that nests most existing task-based models of technical change.
@@ 1  Introduction
Generative artificial intelligence (GenAI) based on large language models (Vaswani et al., 2017) is not a single “machine.” It is a directional technology: it excels at some task combinations and performs poorly at others. Like other general-purpose technologies (Eloundou et al., 2023), it may reshape broad swaths of work, but two features make adoption and gains unusually hard to anticipate ex ante. First, adoption in knowledge work is often worker-driven: the interface is general-purpose and the main inputs are time and attention rather than specialized capital, so usage can spread inside organizations without explicit firm-level decisions. [^1] On worker-driven adoption (”shadow IT”), see Klotz et al. (2019); Raković et al. (2020); for rapid early diffusion, see UBS (2023). Second, capabilities improve quickly but unevenly across tasks. [^2] For software engineering benchmarks, see SWE-bench (Jimenez et al., 2024). The unevenness is described as a ”jagged technological frontier” by Dell’Acqua et al. (2023) and quantified more broadly by Hendrycks et al. (2025). Together, worker-driven adoption and jagged capability mean that asking “what is the effect of AI on labor?” is ill-posed without specifying which tools help with which tasks—and for which workers.
Adoption is especially hard to characterize in knowledge work because workers can also change how they do the job. Job requirements may be stringent in terms of expected output, but different task mixes can deliver the same valued output. A software engineer can allocate time between writing new code, debugging, documentation, and code review; a management consultant between data analysis, client meetings, and slide preparation; a lawyer between legal research, drafting, and client communication. Reallocating effort across tasks is costly, however, and using AI consumes the same scarce time and attention that workers devote to their own work.
To understand AI adoption incentives in knowledge work, we develop a model in which what a worker can do in a unit of time is a convex set of feasible task combinations—a production possibility set (PPS). [^3] This move is standard in international trade, where convex feasible sets and their supporting prices summarize production opportunities (Dixit and Norman, 1980). A worker is employed in a job, which determines the value of any bundle of tasks (Autor, 2013, 2025; Lise and Postel-Vinay, 2020). [^4] For evidence that workers have multidimensional skill endowments that affect occupational mobility, see Gathmann and Schoenberg (2010). They choose a task mix to maximize output subject to a time budget. Skills determine which task combinations are feasible; job requirements determine which combinations generate high output.
We then treat AI as another production possibility set: a potential coworker with its own feasibility profile, and we model it as a directional technology that produces tasks in a fixed proportion. AI capability determines how much it can produce per unit of supervision time, in that fixed proportion. Using AI requires supervision, which uses time the worker could otherwise spend on their own tasks. The combined feasible set is therefore a weighted sum of what the worker and the tool can each produce, where the weights reflect how the worker allocates time between doing and supervising. Adopting AI is a collaboration decision rather than a simple substitution decision: the worker adopts when the expanded set is valuable enough to justify supervision. [^5] The same logic extends to tools that can flexibly combine a small number of task directions in a high-dimensional space; see Kaushik et al. (2025).
The worker adopts an AI tool if using it raises their output. At their pre-AI optimum, each task direction has a shadow (autarky) price: the output gain from being able to do slightly more of that task, holding everything else fixed. These shadow prices summarize what is scarce in the worker–job match and map the AI tool’s direction—a vector in task space—into a single scalar value. The worker adopts the AI tool when its direction is valuable at autarky prices. Autarky prices need not match the worker’s activity vector: they reflect what they would like to shift toward at the margin, not what they currently do.
As AI capability improves, the set of adopted directions expands in a cone centered on the worker’s autarky prices; near the entry threshold, small capability gains translate into large extensive-margin expansions in adoption. Adoption also has a structured intensive margin: a tool can be worth using but not worth using all the time, generating a region of stable partial adoption between an entry threshold and an all-in threshold. Because autarky prices depend jointly on worker skills and job requirements, adoption is match-specific: neither “which workers adopt” nor “which jobs are affected” has a well-defined answer without specifying the pairing. Finally, greater flexibility—in how the worker can allocate their time or in how different tasks are valued in output—widens the set of tool directions that are worth adopting, so adoption expands faster as capabilities improve.
@@
We keep the analysis intentionally neoclassical to make the mechanism transparent. We study a worker doing a given job and abstract from governance and principal-agent frictions, so the worker maximizes output. The main objects and results—autarky prices, the adoption condition, the adoption cone, and the intensive margin—do not depend on a particular parametric form; they follow from convexity and marginal valuation. [^6] Formally, we characterize task choice using a trade-style duality argument that delivers supporting prices for a convex feasible set (Dixit and Norman, 1980).
For measurement and interpretation, we also provide a parametric illustration using standard constant-elasticity functional forms. We assume a constant-elasticity-of-substitution structure for how tasks combine in output, and a constant-elasticity-of-transformation (CET) structure (Powell and Gruen, 1968) for how a worker’s fixed time/attention budget can be used for different task bundles. This specialization yields closed-form expressions for autarky prices as functions of observable skill and requirement vectors and clarifies when skills versus job requirements dominate adoption predictions. In the limiting cases, rigid specialists have shadow prices pinned down by job requirements, while perfectly flexible workers have shadow prices pinned down by skills. In between, which we believe is the relevant case for knowledge work, both matter.
We hold the worker–job match fixed and analyze adoption of a given tool as an individual optimization problem. This isolates a building block that can be embedded in richer settings. In equilibrium, the availability of AI can reshape sorting and the distribution of jobs; multiple tools raise a team-composition problem; firms may govern usage and provide complementary investments (Aldasoro et al., 2026); and tool directions and capabilities may themselves respond to training data and market incentives.
@@
The remainder of the paper proceeds as follows. Section 2 develops the high-dimensional model of task choice and directional technology adoption and derives the adoption condition, the adoption cone, and the intensive margin. Section 3 provides the parametric specification and its closed-form expressions. Section 4 relates the framework to prior task-based models and discusses interpretation and limitations. The Appendix collects proofs and additional derivations.
@@ 2  A Model of Digital Technology Adoption
We study a worker doing a given job. We abstract from labor-market, governance, and principal-agent frictions, so the worker maximizes output. This corresponds to a richer environment where tasks are observable and contractible within the firm. We take the worker–job match as fixed and reserve the study of sorting for future work.
The worker chooses how much to do of each of $N$ tasks. Denote the task quantities by the vector $x\in\mathbb{R}_{+}^{N}$. The set of task bundles the worker can complete in a workday (or other unit of the scarce resource) is given by the set $X_{A}$, characterized by its resource use function $g$,
@@
Here $B$ is the amount of the scarce resource (“budget”) the worker has available. We interpret the budget as time or attention, not money. This is a key feature of digital technology adoption: the binding constraint is often the worker’s cognitive capacity and hours, not financial cost. The set $X_{A}$ is called the Production Possibilities Set (PPS).
@@ 2.1  Technology expands the production possibilities set
To use AI, the worker has to use some of the same resources that constrain them in their own work. For example, instructing AI agents and reviewing their work takes time, and this limits the amount of AI that can be used in a workday. [^8] Time as a binding constraint is central to the economics of digital goods. Goodman et al. (2026) leverage Becker’s time allocation theory to study consumer demand for time-intensive platforms like Facebook and Instagram, finding that time shares and time costs significantly influence substitution patterns across online activities. We assume AI is competing for the same resource budget $B$, so if the worker spends a fraction $\lambda$ of their budget supervising AI, only $(1-\lambda)B$ is left for doing other tasks. [^9] For simplicity, we assume there is a linear transformation between resources spent on AI supervision and resources spent on human work. This is eminently true if the resource is time: one hour lost on an activity is one hour gained on another. A more general formulation could model the production possibilities set as constrained by several resources, such as time, attention, mental energy, and AI may not use them in the same proportion as human tasks. We conjecture our results would survive as long as AI is sufficiently substitutable to human tasks in the resource constraint. We denote the capability of technology with $\chi\geq 0$, which acts as a productivity shifter. A unit of technology $\|t\|$ costs $1/\chi$ units of the resource $B$. Continuing with the AI example, more capable AI tools can get more done per hour of human supervision.
@@
Absolute advantage is necessary but not sufficient for technology adoption. A counterexample is illustrated in Figure 4. The technology already has an absolute advantage, so the worker could use it to expand their PPS. In their current job, however, their activities are directionally far from the technology, so it is not worth adopting it. [^11] Geometrically, until the technology vector pierces the budget set defined by the hyperplane separating the PPS from the production function upper contour set, the convex hull of the technology point and the PPS remains strictly within the budget set and offers no revenue-increasing opportunities to the worker.
@@ 2.3  The intensive margin of technology
Suppose $F$ is strictly quasi-concave and $C^{2}$. Let $y=B\chi t$ and $X_{T}=\mathrm{conv}(X_{A}\cup\{y\})$. Then:
@@
Uniqueness of $\tilde{x}_{*}$ follows from strict quasi-concavity of $F$ over the convex set $X_{T}$. The decomposition follows from the convex hull representation (Proposition 2.3). Uniqueness of $\lambda_{*}$ follows from strict concavity of $f(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$ (which holds because $F$ is strictly quasi-concave and HD1), and $x_{*}$ is then pinned down by $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$. Uniqueness of $p_{*}$ follows from $C^{2}$ differentiability of $F$. See Appendix A.1. ∎
@@
Suppose $F$ is strictly quasi-concave, $C^{2}$, and homogeneous of degree one. Let $\lambda_{*}(\chi)$ denote the unique optimal technology intensity from Proposition 2.7. Then:
@@
Cases 1–2 follow from the boundary conditions of the univariate function $f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]$. For case 3, the strict inequality $\chi_{100}>\chi_{0}$ follows from strict quasi-concavity: the supporting hyperplanes at $x_{A}$ and at the technology direction $t$ are distinct whenever $t$ is not collinear with $x_{A}$. See Appendix A.2. ∎
@@ 2.4  More than one technology
    All-in: The worker goes all-in on technology $K+1$ if $\chi_{K+1}>\varrho_{K}(p_{100,K+1})/p_{100,K+1}^{\prime}t_{K+1}$ and $\chi_{K+1}/\chi_{k}>(p_{100,K+1}^{\prime}t_{k})/(p_{100,K+1}^{\prime}t_{K+1})$ for all $k=1,\ldots,K$, where $p_{100,K+1}=\nabla F(t_{K+1})/\|\nabla F(t_{K+1})\|$.
@@ 3.1  CES job requirements and CET skills
Skills (CET). The resource use function is constant elasticity of transformation (CET) (Powell and Gruen, 1968):
@@ 3.2  Autarky allocation in closed form
$x_{i}^{A}=B\cdot\frac{s_{i}^{\sigma/(\gamma+\sigma)}\,\theta_{i}^{\gamma/(\gamma+\sigma)}}{\Phi^{\gamma/(\gamma+1)}},\qquad i=1,\ldots,N,$
@@
The allocation formula shows that the worker concentrates activity where skills and requirements overlap. The exponents make the role of curvature transparent. When $\gamma$ is high, the worker can freely reallocate across tasks, so they tilt toward what the job requires (higher weight on $\theta_{i}$). When $\sigma$ is high, tasks are easily substitutable in production and requirements matter less, so the worker leans toward tasks where they are skilled (higher weight on $s_{i}$).